### v2

# Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study over Open-ended Question Answering

Abstract
Recent works integrating Knowledge Graphs (KGs) have led to promising improvements in enhancing the reasoning accuracy of Large Language Models (LLMs). However, current benchmarks focus mainly on closed-ended tasks, leaving a gap in the assessment of more complex real-world scenarios. This gap has also obscured the evaluation of KGs’ potential to mitigate the problem of hallucination in LLMs. To fill the gap, we introduce OKGQA, a new benchmark specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions from different types, and incorporates specific metrics to measure both hallucination ratio and the enhancement in reasoning capabilities. To consider the scenario in which KGs may have varying levels of mistakes, we propose another benchmark variant OKGQA-P to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated. OKGQA aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design. We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination. The code and data are released at link.
Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study over Open-ended Question Answering
Yuan Sui1, Yufei He1, Zifeng Ding2, Bryan Hooi1 1 National University of Singapore 2 University of Cambridge {yuansui, yufei.he, bhooi}@comp.nus.edu.sg, zd320@cam.ac.uk
1 Introduction
Contemporary LLMs are prone to producing hallucinations due to gaps in their knowledge (Gekhman et al., 2024; Lee et al., 2023), primarily when the training data contain misinformation, biases, or inaccuracies. These flaws can arise when we simply maximize the log-likelihood, leading to responses that may seem plausible, but are often irrelevant or incorrect (Weng, 2024). This issue is especially concerning in scenarios where accuracy and reliability are essential, such as healthcare (He et al., 2023) and science (Taylor et al., 2022).
Figure 1: Comparison between Open-ended Question and Close-ended Question over Knowledge Graph.
To address this critical limitation, researchers have turned to leveraging external knowledge graphs (KGs) as a complementary (Pan et al., 2024; Luo et al., 2023a; Hu et al., 2023; Sun et al., 2023; van Breugel et al., 2021). KGs provide several key advantages: they contain structured and explicit factual knowledge, including domain-specific information, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023). Furthermore, KGs offer traceability, allowing each piece of information to be traced back to its source. This traceability enables both verification of generated context and transparency in reasoning paths, making KGs a promising source for improving the reliability of LLM reasoning. Find more details of related works in §B.
However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities or relations (Talmor et al., 2019; Mihaylov et al., 2018; Jin et al., 2020) or a set of logical forms (Yih et al., 2016; Talmor and Berant, 2018; Puerto et al., 2023) derived from specific facts of KG (as shown in Figure 1). These benchmarks, while useful for evaluating basic retrieval and reasoning, are insufficient for testing the LLMs’ ability for faithful generation (i.e., minimize hallucination in LLM responses), leaving a gap in the assessment of complex, real-world scenarios.
Why Open-ended Question? A key limitation of closed-ended benchmarks is: they cannot assess the phenomenon of “more talk, more mistakes”, where longer, more complex responses tend to increase hallucination Qiu et al. (2024). Unlike closed-ended benchmarks that require short, specific answers, open-ended questions demand more detailed and nuanced explanations. This format also enables direct evaluation metrics like FActScore (Min et al., 2023) and SAFE (Wei et al., 2024), which measure factual consistency of LLM responses by decomposing them into atomic statements for comparison with external information. However, for short, rigid responses, this decomposition is not feasible, rendering these metrics inapplicable. Moreover, open-ended evaluations align more closely with human cognition Lu et al. (2024); Amirizaniani et al. (2024), where context-rich and interpretative answers are expected compared to simplistic and mechanical responses. Find more details of intuition in Appendix §A.1.
Motivation: Recognizing the research gap, our study aims to contribute to the existing knowledge from two perspectives: 1) explore whether KGs can make LLMs more trustworthy in the open-ended setting, and 2) conduct a comparative analysis to shed light on methods and directions for leveraging KGs to reduce LLMs’ hallucination.
To achieve this, we develop a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs under open-ended question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions of different types and incorporates specific metrics to measure both the reduction in hallucinations and enhancements in reasoning capabilities. Specifically, all queries in OKGQA are synthesized by LLMs with pre-defined templates mentioned in Table 1 and corresponding Wikipedia context. To consider the scenarios in which KGs may have varying levels of mistakes (i.e., attributes may be mislabeled, relations may not exist, etc.), we propose another experiment setup OKGQA-P (§2.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. We assess the hallucination ratio and the quality of the response (detailed metrics can be found in §4.1) of the tested models.
Based on our experiments, we find that (1) retrieving KG information can mitigate factual errors in LLMs’ responses, especially for queries requiring additional reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thoughts (Kim et al., 2023) and self-consistency (Wang et al., 2022)) may cause bias and hallucination; (3) Subgraphs generally achieve the best performance across different query types, especially for simpler types of queries; (4) Integrating KGs can effectively reduce hallucinations in LLMs, even when KGs are contaminated to a certain degree.
2 OKGQA: A Open-ended Knowledge Graph Question-Answering Benchmark
Overview. OKGQA is a comprehensive benchmark designed to assess how effectively LLMs enhanced with KGs perform in open-ended, real-world-like question answering scenarios. Unlike existing benchmarks that focus primarily on closed-ended tasks, OKGQA presents diverse open-ended question types that mirror the variable nature of practical applications. The benchmark evaluates how effectively models leverage KGs to generate both reliable and comprehensive answers, focusing on reducing hallucinations while maintaining reasoning capabilities of LLMs. To assess robustness under imperfect conditions, we also introduce OKGQA-P, a variant where KGs are deliberately perturbed and contaminated.
2.1 Dataset Construction
Statistics (on average)
Tokens in query 23.97
Total number of queries 850
Number of unique DBPedia entities 816
Before Pruning $\rightarrow$ After PPR Pruning
Tokens in subgraph 348,715 $\rightarrow$ 2,452
Number of nodes 7,171 $\rightarrow$ 48
Number of Edges 8,213 $\rightarrow$ 152
Avg. Degree 1.15 $\rightarrow$ 3.17
Clustering Coefficient 0.00 $\rightarrow$ 0.69
Graph Density 0.00 $\rightarrow$ 0.07
Query Type Simple Complex
Descriptive 78 11
Explanatory 195 56
Predictive 110 55
Comparative 72 74
Critical 182 17
Total 637 213
(a) Dataset statistics and query types
(b) Distribution of sub-query types
Figure 2: (left) Dataset statistics and query types, (right) Sub-query type distribution
Queries. We utilize a template-based methodology to generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries. Details regarding specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix C. To ensure that the generated queries represent real-world scenarios and complexities, we employ an iterative optimization approach that utilizes both automated and human evaluation to refine the query generation process (the details are given in Appendix A.2). Initially, we generate a diverse set of query candidates from a seed instruction. These candidates undergo automated evaluation using an LLM-based evaluator, which assigns quality scores $s_{\text{auto}}$ on a scale of 1-10, with higher scores indicating better performance across multiple metrics. Subsequently, human evaluators assess the same queries, producing corresponding normalized scores $s_{\text{human}}$ within the same range. To refine the dataset, we iteratively optimize the input instructions by minimizing the discrepancy between $s_{\text{human}}$ and $s_{\text{auto}}$. This optimization process ensures alignment between automated and human quality assessments. The queries are also categorized by complexity with detailed statistics in Figure 2.
KG Sub-graphs. To reduce the size of KGs, following previous works (Yih et al., 2016; Talmor and Berant, 2018), we sample subgraphs from DBpedia Knowledge Graph (Noted that all the queries in OKGQA can be reasoned based on DBPedia) by extracting all triples contained within the k-hop neighbors from the question entities in the query. We set k = 2 to balance graph size and computational feasibility, as increasing k leads to exponential growth in the number of edges and nodes, which can introduce excessive noise and make it more challenging to retrieve essential information from KGs [^1] This choice is also informed by common practices in other benchmarks, such as WebQSP Yih et al. (2016) and CWQ Talmor and Berant (2018), where 2-hop subgraphs are widely used for similar KGQA tasks.. To further reduce the size of the 2-hop subgraphs, we leverage personalized page-rank (PPR) (Bahmani et al., 2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.3). We compare the statistics of subgraphs before and after PPR pruning in Figure 2(a).
2.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
To further mimic the real situations where KGs may not be of high quality (i.e., attributes of nodes/edges may be mislabeled, relations may not exist, etc.), we propose another experiment setting OKGQA-P to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. We introduce various perturbation scenarios including mislabeled attributes, incorrect relations, and missing connections to test the model’s reasoning capabilities with flawed or incomplete KG data. To quantify the degree of perturbation, we also evaluate both the semantic and structural similarity between the original and the modified KG as defined below.
Notation. Let $\mathcal{F}_{\theta}$ be a KG-augmented model, and KG as $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{T})$, where $\mathcal{V}$ is the set of entities (nodes), $\mathcal{E}$ is the set of relation types (edges), and $\mathcal{T}=\{(v_{1},e,v_{2})|v_{1},v_{2}\in\mathcal{V},e\in\mathcal{E}\}$ is the set of triplets composed of entities and relations. Let $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E}^{\prime},\mathcal{T}^{\prime})$ be the KG after perturbing $\mathcal{G}$, where $\mathcal{E}^{\prime}\neq\mathcal{E}$ and $\mathcal{T}^{\prime}\neq\mathcal{T}$. Let $f(\mathcal{G},\mathcal{G}^{\prime})$ be a function that measures the similarity between $\mathcal{G}$ and $\mathcal{G}^{\prime}$. Let $g(\mathcal{G})$ be the downstream performance when evaluating $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$.
High-level Procedure. First, we test $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}$ to get the original performance $g(\mathcal{G})$. Second, we perturb $\mathcal{G}$ to obtain $\mathcal{G}^{\prime}$. Third, we evaluate $\mathcal{F}_{\theta}$ on data samples $X$ and $\mathcal{G}^{\prime}$ to get the perturbed performance $g(\mathcal{G}^{\prime})$. Finally, we measure $g(\mathcal{G})-g(\mathcal{G}^{\prime})$ and $f(\mathcal{G},\mathcal{G}^{\prime})$ to assess how robust $\mathcal{F}_{\theta}$ is, i.e., to assess the model performance under conditions where KGs’ semantics and structure are deliberately perturbed.
To quantify how much the perturbed KG has deviated from the original KG, i.e., $f(\mathcal{G},\mathcal{G}^{\prime})$, we leverage metrics from Raman et al. (2020) for capturing semantics (ATS) and structural (SC2D, SD2) similarity between KGs. Intuitively, ATS leverages a pre-trained LM for link prediction to measure the probability of each edge from $\mathcal{G}^{\prime}$ existing in $\mathcal{G}$, while SC2D and SD2 measure the structural similarity between two KGs based on local clustering coefficient and degree distribution. For each of the three metrics, higher value indicates higher similarity. The detailed description can be found in Appendix A.5, with visualization in Figure 5.
For the perturbation methods, we consider four perturbation heuristics based on (Raman et al., 2020) as follows: Relation Swapping (RS) randomly chooses two edges from $\mathcal{T}$ and swaps their relations. Relation Replacement (RR) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, and replaces the $e_{1}$ with another relation $e_{2}=\text{argmin}_{e\in\mathcal{E}}S_{\mathcal{G}}(v_{1},e,v_{2})$, where $S_{\mathcal{G}}(\cdot)$ is a KG score function adapted from ATS. The replacement is performed such that the resulting $(v_{1},e_{2},v_{2})$ has the closest score to the original triplet $(v_{1},e_{1},v_{2})$ to generate “harder negatives” - triplets that are semantically similar but incorrect, thereby challenging the model to different subtle distinctions. Noted that $S_{G}$ differs from the original ATS, as it operates on individual triplets rather than comparing two entire KGs. Edge Rewiring (ER) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$, then replaces $v_{2}$ with another entity $v_{3}\in\mathcal{E}\backslash\mathcal{N}_{1}(v_{1})$, where $\mathcal{N}_{1}(v_{1})$ represents the 1-hop neighborhood of $v_{1}$. Edge Deletion (ED) randomly chooses an edge $(v_{1},e,v_{2})\in\mathcal{T}$ and deletes it. We control the perturbation level based on the percentage of KG edges perturbed. See results in Figures 5 and 6.
Table 1: Query types and examples in OKGQA. Brown is used to highlight the placeholders (e.g., [person], [event]) in description, while Teal highlights the specific entities to replace the placeholders.
Type Sub-Type Description / Template Example
Descriptive Character Description
Describe a [person]’s significant contributions
during their career.
Please describe Albert Einstein’s contributions to the
field of physics.
Event Description
Provide a detailed description of the background
and course of an [event].
Please provide a detailed description of the background
and course of the French Revolution.
Explanatory Cause Explanation
Why did [person] take [action] at [time]?
Why did Nixon choose to resign from the presidency
in 1974?
Relationship Explanation
Explain the relationship between [entity A] and
[entity B] and its significance.
Explain the relationship between Alexander the Great
and Aristotle and its significance.
Predictive Trend Prediction
Based on the historical behavior of [entity], what
do you think it might do in the future?
Based on Tesla’s historical behavior, in which fields do
you think it might innovate in the future?
Outcome Prediction
Based on the current situation, how do you predict
[event] will develop?
Based on the current international situation, how do you
predict climate change policies will develop?
Comparative Contrast Analysis
Compare and contrast the similarities and differences
between [entity A] and [entity B] in [aspect].
Compare and contrast the leadership styles of Steve Jobs
and Bill Gates.
Historical Comparison
Compare the impact of [historical event A] and
[historical event B].
Compare the impact of World War I and World War II
on the global order.
Critical Evaluation and Reflection
How do you evaluate the impact of [person/event]
on [field]? Please explain your viewpoint.
How do you evaluate Martin Luther King’s impact on
the civil rights movement? Please explain your viewpoint.
Application and Practice
How do you think [theory/method] can be applied
to [practical issue]?
How do you think machine learning technology can be
applied to medical diagnostics?
3 Exploring KG-augmented framework for reducing hallucination
To explore whether KG-augmented approaches can mitigate LLMs’ hallucination, we investigate a range of methods and discover that they can be encapsulated within a unified framework as shown in Figure 3. Our framework follows the paradigm of retrieval augmented generation (RAG) (Edge et al., 2024; Baek et al., 2023a), which retrieves essential information from the KGs, and then uses the retrieved knowledge to enhance the LLM’s generation. It has two components, i.e., Graph-guided retrieval (G-retrieval) and Graph-guided generation (G-generator), with a variety of algorithmic design choices. We analyze our proposed strategies within each component in §4, aiming to shed light on the best practices for leveraging KGs for reducing hallucinations in LLMs.
Figure 3: Overview of KG-augmented framework.
Here we formalize the problem of context-relevant KG retrieval for open-ended question answering. Given a user query $q$, a pretrained language model (PLM) first encodes the input tokens, and models a conditional distribution $p(a|q)$ to generate an output response $a=[a_{1},\ldots,a_{T}]$. To explore whether KGs help reduce hallucinations of LLMs, we introduce the retrieved knowledge from the KG, denoted $\mathcal{Z}\in\mathcal{G}$, which can take various forms (e.g., triplets, paths, subgraphs, as shown in Figure 3). This retrieved knowledge contains information relevant to the user query, and either benefits the LLMs at its knowledge boundaries (Baek et al., 2023b) or explicitly guides its reasoning process (Luo et al., 2023a; Sun et al., 2023). The distribution of the retrieved knowledge $\mathcal{Z}$ is modeled as $p(\mathcal{Z}|q,\mathcal{G})$, leading to a new likelihood for generating responses, expressed as $p(a|q)$.
$p(a|q)$ $=\sum_{\mathcal{Z}\subseteq\mathcal{G}}p_{\phi}(a|q,\mathcal{Z})p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ (1)
$\approx p_{\phi}(a|q,\mathcal{Z}^{*})p_{\theta}(\mathcal{Z}^{*}|q,\mathcal{G}),$
where $\mathcal{Z}^{*}=\text{argmax}_{\mathcal{Z}\in\mathcal{G}}p_{\theta}(\mathcal{Z}|q,\mathcal{G})$ represents the optimal subset of retrieved knowledge from the KGs, and $p_{\phi}(a|q,\mathcal{Z}^{*})$ represents the likelihood generating the response given the query $q$ and the retrieved knowledge $\mathcal{Z}^{*}$, parameterized by $\phi$ and $\theta$. Due to the potential exponential growth in the number of possible retrieved knowledge with graph size, efficient approximation methods are required. Consequently, the first line of Eq. 1 is approximated using the most probable retrieved knowledge $\mathcal{Z}^{*}$.
3.1 Graph-guided retrieval (G-retrieval)
In this section, we explore which forms of knowledge (e.g., triplets, paths, or subgraphs) should be extracted from KGs to enhance the LLMs and discuss the design of G-retrieval. The phase begins by encoding the user query into an embedding vector $\mathbf{q}\in\mathbb{R}^{d}$ using a language model encoder $f_{\text{LM}}(\cdot)$. For each node $v\in\mathcal{V}$ and edge $e\in\mathcal{E}$, their respective embeddings $\mathbf{e}_{v}$ and $\mathbf{e}_{e}$ are similarly computed using the same model, ensuring all the information is represented in a unified vector space. The relevance of each node or edge to the query is measured using cosine similarity as: $s_{v}=\frac{\mathbf{q}\cdot\mathbf{e}_{v}}{\|\mathbf{q}\|\|\mathbf{e}_{v}\|}$ and $s_{e}=\frac{\mathbf{q}\cdot\mathbf{e}_{e}}{\|\mathbf{q}\|\|\mathbf{e}_{e}\|}$. Based on these similarity scores, the most relevant nodes and edges are identified to form a set of top-$k$ nodes and edges for the query.
To balance retrieving as many relevant nodes and edges as possible while keeping the size manageable, we adopt a prize-cost trade-off strategy: (1) Prize assignment: based on the computed similarity scores, we assign prizes to nodes and edges to quantify their relevance to the query. Specifically, we assign the top-$k$ nodes/edges with descending prize values from $k$ to 1, while nodes and edges outside the top-$k$ receive a prize of 0. Formally: $p_{v}=\max(0,k-\text{rank}(v)+1)$ and $p_{e}=\max(0,k-\text{rank}(e)+1)$. (2) Cost allocation: to manage the retrieved knowledge size, we assign penalties as cost $C_{e}$ during the expansion of the retrieved paths or subgraphs. These prizes and costs will be used in the retrieval process, by aiming to maximize the prizes while minimizing costs.
Triplet-retrieval: the triplet retrieval approach retrieves a fixed number $k$ of triplets with the highest total prize assigned to their respective head entity, relation, and tail entities.
Path-retrieval: based on the defined prize assignment and cost allocation, the path-retrieval approach starts from a fixed number of $k$ of high-prize nodes to construct sequences $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$ by aiming to greedily maximize their scores: $S(\mathcal{P})=\sum_{i=1}^{n}p_{v_{i}}+\sum_{i=1}^{n-1}p_{e_{i}}-\sum_{i=1}^{n-1}c_{e}$. We use a priority queue to iteratively extend paths by selecting the highest-scoring extensions while avoiding cycles and adhering to maximum path length and quantity constraints. Finally, the top-scoring paths are sorted and returned. The details of path-retrieval can be found in Appendix A.4.
Sub-graph retrieval: the sub-graph retrieval approach also uses the prize assignment and cost allocation, following He et al. (2024a) to optimize the process using Prize-Collecting Steiner Tree (PCST) algorithm. The PCST problem seeks to build a connected subgraph $\mathcal{S}$ that maximizes the total prizes of included nodes and edges while minimizing the total cost of the edges. The score of the subgraph is defined as $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{einE_{\mathcal{S}}}c_{e}$. Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
3.2 Graph-guided generation (G-generator)
After retrieving $\mathcal{Z}^{*}$, the graph-guided generation phase learns to use this knowledge to generate the corresponding output to answer the user query. The generation is modeled as a sequential decision-making process, where at each time step $t$, the next token $a_{t}$ is generated by G-Generator, conditioned on the query $q$, the retrieved knowledge $\mathcal{Z}^{*}$, and all previously generated tokens $a_{0:t-1}$.
$p(a|q,\mathcal{Z}^{*})=\prod_{t=1}^{T}p_{\theta}\left(a_{t}|q,\mathcal{Z}^{*},a_{0:t-1}\right)\vspace{-2mm}$ (2)
The model learns the parameters $\theta$ to maximize the likelihood of generating the correct output sequence by computing the probability distribution over possible next tokens at each time step. This ensures that the generated output not only considers the previously generated tokens but also the retrieved knowledge. The generation continues until a termination condition is met, such as generating an end-of-sequence token or reaching the maximum sequence length $T$.
4 Experiments
4.1 Evaluation Metrics
To quantify the hallucinations of LLMs, we leverage two public metrics, FActScore (Min et al., 2023) (factual precision in atomicity score) and SAFE (Wei et al., 2024) (search-augmented factuality evaluator). FActScore is designed to measure factual precision in text by decomposing a long form generation into multiple atomic facts and validates each separately against a reliable knowledge base, such as Wikipedia. We measure the proportion of facts that are supported by the knowledge source out of the total atomic facts. SAFE, on the other hand, adopts a more dynamic approach to fact-checking. It employs a language model as an investigative agent, which iteratively uses Google Search queries and reasons about whether the search results support or do not support the fact.
Figure 4: Comparison results of different forms of information over different queries.
Table 2: Comparison results of various forms of information extracted from the KGs.
Models G-Eval Hallucination
Context Relevance Comprehensiveness Correctness Empowerment SAFE FActScore
Baseline: Without External Knowledge (Zero-shot prompting)
GPT-4o $68.12\%\pm 0.88\%$ $65.41\%\pm 0.79\%$ $60.41\%\pm 0.38\%$ $62.41\%\pm 0.84\%$ $82.47\%\pm 0.62\%$ $55.34\%\pm 0.93\%$
GPT-4o-mini $63.21\%\pm 0.49\%$ $60.11\%\pm 0.47\%$ $55.43\%\pm 0.63\%$ $58.72\%\pm 0.62\%$ $80.14\%\pm 0.89\%$ $50.23\%\pm 1.01\%$
llama-3.1-8b-instruct $57.12\%\pm 0.91\%$ $54.74\%\pm 1.20\%$ $49.01\%\pm 0.61\%$ $52.21\%\pm 0.71\%$ $79.33\%\pm 0.91\%$ $45.14\%\pm 0.32\%$
mistral-7B-Instruct-v0.3 $55.71\%\pm 1.21\%$ $52.00\%\pm 1.31\%$ $47.03\%\pm 0.94\%$ $50.13\%\pm 1.04\%$ $78.27\%\pm 0.83\%$ $44.37\%\pm 1.23\%$
gemma-2-9b-it $53.63\%\pm 1.33\%$ $50.00\%\pm 1.33\%$ $45.72\%\pm 0.71\%$ $48.15\%\pm 0.93\%$ $77.11\%\pm 0.78\%$ $40.94\%\pm 0.83\%$
Baseline: Without External Knowledge (4-shot prompting)
GPT-4o $70.61\%\pm 0.62\%$ $67.43\%\pm 0.81\%$ $62.33\%\pm 0.37\%$ $64.51\%\pm 0.12\%$ $83.39\%\pm 0.53\%$ $57.45\%\pm 0.78\%$
GPT-4o-mini $65.53\%\pm 0.94\%$ $62.33\%\pm 1.03\%$ $57.23\%\pm 0.68\%$ $60.47\%\pm 0.83\%$ $81.62\%\pm 0.69\%$ $52.34\%\pm 0.76\%$
llama-3.1-8b-instruct $59.43\%\pm 0.32\%$ $56.31\%\pm 0.78\%$ $51.27\%\pm 0.32\%$ $54.33\%\pm 0.41\%$ $80.27\%\pm 0.78\%$ $47.24\%\pm 1.03\%$
mistral-7B-Instruct-v0.3 $57.34\%\pm 1.04\%$ $54.13\%\pm 1.31\%$ $49.27\%\pm 0.84\%$ $52.46\%\pm 0.94\%$ $79.12\%\pm 0.87\%$ $45.13\%\pm 1.42\%$
gemma-2-9b-it $55.24\%\pm 1.49\%$ $52.27\%\pm 1.21\%$ $47.14\%\pm 0.36\%$ $50.36\%\pm 0.51\%$ $78.00\%\pm 0.77\%$ $44.32\%\pm 1.58\%$
Var-1: With CoT Prompting
GPT-4o - CoT $72.76\%\pm 0.92\%$ $69.56\%\pm 0.74\%$ $64.48\%\pm 0.63\%$ $66.69\%\pm 0.69\%$ $80.07\%\pm 0.83\%$ $54.30\%\pm 0.87\%$
GPT-4o - CoT+SC $75.81\%\pm 0.65\%$ $71.62\%\pm 0.74\%$ $66.55\%\pm 0.75\%$ $68.74\%\pm 0.15\%$ $79.03\%\pm 0.48\%$ $53.23\%\pm 0.78\%$
llama-3.1-8b-instruct - CoT $61.54\%\pm 0.95\%$ $58.35\%\pm 1.05\%$ $53.31\%\pm 0.71\%$ $56.42\%\pm 0.83\%$ $77.07\%\pm 0.85\%$ $46.15\%\pm 0.54\%$
llama-3.1-8b-instruct - CoT+SC $63.69\%\pm 0.32\%$ $60.44\%\pm 0.59\%$ $55.46\%\pm 0.52\%$ $58.53\%\pm 1.11\%$ $76.00\%\pm 0.63\%$ $45.05\%\pm 0.97\%$
mistral-7B-Instruct-v0.3 - CoT $59.58\%\pm 0.43\%$ $56.23\%\pm 2.31\%$ $51.28\%\pm 1.31\%$ $54.33\%\pm 0.72\%$ $75.04\%\pm 0.95\%$ $43.03\%\pm 1.03\%$
mistral-7B-Instruct-v0.3 - CoT+SC $61.35\%\pm 0.93\%$ $58.33\%\pm 1.02\%$ $53.42\%\pm 0.79\%$ $56.47\%\pm 0.85\%$ $74.30\%\pm 0.21\%$ $42.00\%\pm 0.29\%$
gemma-2-9b-it - CoT $57.34\%\pm 1.05\%$ $54.12\%\pm 0.32\%$ $49.27\%\pm 0.85\%$ $52.12\%\pm 0.95\%$ $72.07\%\pm 1.05\%$ $40.13\%\pm 0.49\%$
gemma-2-9b-it - CoT+SC $59.42\%\pm 0.27\%$ $56.27\%\pm 0.84\%$ $51.34\%\pm 1.42\%$ $54.34\%\pm 1.31\%$ $71.09\%\pm 0.43\%$ $39.85\%\pm 1.03\%$
Var-2: With Triplets Extracted from KGs Provided
GPT-4o $74.62\%\pm 0.65\%$ $70.44\%\pm 0.79\%$ $65.37\%\pm 0.72\%$ $67.12\%\pm 0.71\%$ $89.20\%\pm 1.42\%$ $72.53\%\pm 0.83\%$
GPT-4o-mini $69.50\%\pm 0.81\%$ $65.03\%\pm 0.92\%$ $60.21\%\pm 0.65\%$ $63.43\%\pm 1.01\%$ $87.52\%\pm 0.34\%$ $67.73\%\pm 0.95\%$
llama-3.1-8b-instruct $63.45\%\pm 1.13\%$ $59.33\%\pm 1.05\%$ $54.23\%\pm 0.75\%$ $57.33\%\pm 0.12\%$ $85.37\%\pm 0.72\%$ $62.37\%\pm 0.82\%$
mistral-7B-Instruct-v0.3 $61.34\%\pm 0.31\%$ $57.21\%\pm 0.89\%$ $52.29\%\pm 0.32\%$ $55.12\%\pm 0.43\%$ $84.21\%\pm 0.84\%$ $60.28\%\pm 1.05\%$
gemma-2-9b-it $59.25\%\pm 1.06\%$ $55.29\%\pm 0.44\%$ $50.15\%\pm 0.85\%$ $53.73\%\pm 0.95\%$ $83.18\%\pm 0.43\%$ $58.13\%\pm 0.91\%$
GPT-4o - CoT+SC $76.71\%\pm 0.53\%$ $72.34\%\pm 0.21\%$ $67.33\%\pm 1.31\%$ $69.64\%\pm 0.33\%$ $88.11\%\pm 0.57\%$ $71.45\%\pm 0.53\%$
Var-3: With Paths Extracted from KGs Provided
[…]
GPT-4o-mini $73.64\%\pm 0.93\%$ $69.41\%\pm 0.22\%$ $64.35\%\pm 0.72\%$ $67.52\%\pm 0.82\%$ $88.22\%\pm 0.34\%$ $70.53\%\pm 0.24\%$
llama-3.1-8b-instruct $67.51\%\pm 0.46\%$ $63.62\%\pm 1.39\%$ $58.41\%\pm 0.93\%$ $61.57\%\pm 0.94\%$ $86.33\%\pm 0.94\%$ $65.42\%\pm 0.95\%$
mistral-7B-Instruct-v0.3 $65.48\%\pm 0.94\%$ $61.37\%\pm 1.01\%$ $56.34\%\pm 0.23\%$ $59.45\%\pm 0.43\%$ $85.26\%\pm 0.85\%$ $63.31\%\pm 1.33\%$
gemma-2-9b-it $63.35\%\pm 1.37\%$ $59.23\%\pm 0.91\%$ $54.31\%\pm 0.91\%$ $57.41\%\pm 0.27\%$ $84.13\%\pm 0.21\%$ $61.23\%\pm 1.04\%$
GPT-4o - CoT+SC $80.87\%\pm 0.42\%$ $76.60\%\pm 0.65\%$ $71.54\%\pm 0.53\%$ $73.79\%\pm 1.21\%$ $89.11\%\pm 0.63\%$ $74.53\%\pm 0.24\%$
Var-4: With Subgraphs Extracted from KGs Provided
[…]
GPT-4o-mini $75.70\%\pm 0.44\%$ $71.51\%\pm 0.83\%$ $66.43\%\pm 0.76\%$ $69.60\%\pm 0.65\%$ $88.71\%\pm 0.72\%$ $70.12\%\pm 0.87\%$
llama-3.1-8b-instruct $69.61\%\pm 0.84\%$ $65.45\%\pm 0.93\%$ $60.41\%\pm 0.65\%$ $63.42\%\pm 0.45\%$ $86.12\%\pm 0.35\%$ $65.44\%\pm 0.87\%$
mistral-7B-Instruct-v0.3 $67.55\%\pm 0.87\%$ $63.35\%\pm 0.43\%$ $58.37\%\pm 0.71\%$ $61.45\%\pm 0.32\%$ $85.21\%\pm 0.81\%$ $63.12\%\pm 0.94\%$
gemma-2-9b-it $65.45\%\pm 0.95\%$ $61.23\%\pm 1.0\%$ $56.31\%\pm 0.35\%$ $59.40\%\pm 0.85\%$ $84.51\%\pm 0.99\%$ $63.74\%\pm 0.49\%$
[…]
In addition to the hallucination metrics, we also propose four metrics to quantify the desirable qualities for generating sensible responses (Edge et al., 2024; Wang et al., 2023). Given the complexities inherent in an open-ended setting, collecting and annotating ground-truth answers poses significant challenges. Therefore, we leverage LLMs as automatic reference-free evaluators to assess these metrics based on G-Eval (Liu et al., 2023), a framework of using LLMs with chain-of-thought (CoT) and a form-filling paradigm. To enhance G-Eval’s robustness and stability, we provide the relevant Wikipedia pages as context in the prompt when calculating the metrics. The metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context. Comprehensiveness: assesses how thoroughly the answer addresses all aspects and details of the question. Correctness: measures the clarity and specificity of the generated answer to the question. Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions. The detailed prompt can be found in Appendix C.
Models. We incorporate a range of widely used LLMs of different scales, including GPT-4o, GPT-4o-mini, Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024). We use the embedding model (text-embedding-3-small) from OpenAI for the G-retrieval process, considering cost and performance.
4.2 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
To explore whether KGs can help reduce hallucination in LLMs, we benchmark the LLMs in different settings. We use zero-shot and few-shot prompting as baselines without injecting external knowledge. In addition, compared to adding external knowledge from KGs, we also consider leveraging LLMs’ internal knowledge to do Chain-of-thought (Kim et al., 2023), or self-consistency (Wang et al., 2022) as baselines. For LLMs augmented with KGs, we consider three KG retrieval forms: triplets, paths, and subgraphs (as shown in Figure 3) to study the impact of G-retrieval for reducing LLMs’ hallucinations. The results are shown in Table 2 and Figure 4. We obtain some intriguing findings:
Retrieving KG information can indeed mitigate factual errors in the responses. Methods integrating knowledge extracted from KGs show clear improvements in factual accuracy and comprehension scores compared to the baselines. For example, under Var-2 (triplet retrieval), GPT-4o achieves a FActScore of 72.55% $\pm$ 0.85%, which is a significant increase over the baseline score of 55.35% $\pm$0.95%. Moreover, these methods can be combined with strategies like CoT+SC, enhancing response quality with minimal increase in hallucination ratio. The radar chart in Figure 4 further emphasizes that in most query types, integrating knowledge retrieved from KGs mitigates the hallucination issue compared to baselines, particularly in query types such as “Evaluation and Reflection,” “Outcome Prediction,” and “Cause Explanation,” which require more reasoning and analysis rather than merely listing information. The findings also apply to open-source models like mistral-7B-Instruct-v0.3 and Llama-3.1-8B-instruct, illustrating the consistency of the finding.
Directly performing reasoning in the LLM itself does not mitigate hallucinations. We benchmark the hallucination ratio of LLMs using internal reasoning strategies like Chain-of-thought and Self-consistency, as shown in Var-1 in Table 2. It shows that these methods can improve response quality (i.e., correctness, comprehensiveness, and empowerment) compared to baselines, but do not consistently improve factuality, and sometimes even diminish factuality. This shows that relying solely on internal reasoning does not mitigate hallucination, emphasizing the importance of external knowledge to counter the hallucination issue.
(a) Edge Deletion
(b) Edge Rewiring
(c) Relation Replacement
(d) Relation Swapping
Figure 5: Performance Metrics (ATS, SC2D, SD2) vs. Perturbation Level for Different Perturbation Methods.
(a) Edge Deletion
(b) Edge Rewiring
(c) Relation Replacement
(d) Relation Swapping
Figure 6: Performance Metric (FActScore) vs. Perturbation Level for Different Perturbation Methods and Different Retrieval Methods. FS-T refers to FActScore metric using triplets, FS-P refers to using paths, and FS-SG refers to using sub-graphs.
Subgraph retrieval generally achieves best performance across different query types, especially for simpler queries. We demonstrate the performance of different retrieval methods across different query types in Figure 4, showing that subgraphs achieve the best performance. Especially for simpler queries (“Character Description” and “Event Description” which do not require intensive reasoning), subgraph retrieval methods achieve better performance compared to paths and triplets. Even for queries like “Relationship Explanation” and “Cause Explanation” which require stepwise reasoning, subgraph methods still demonstrate promising performance. This suggests that while different forms of retrieved knowledge offer unique benefits for specific types of queries, subgraphs provide consistently good performance.
4.3 RQ2: How Are KG-Aware Methods Affected by Noise / Perturbations in KGs?
We benchmark different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality. Specifically, we consider different perturbation methods discussed in §2.2 and control the perturbation level based on the percentage of KG edges being perturbed. We first illustrate how much the perturbed KG has been deviated from the original KG with the increase of perturbation level, shown in Figure 5. It shows that the perturbation methods like edge deletion, rewiring and swapping have relatively weak influence on ATS (which intuitively measures semantic similarity), even as the perturbation level increases. For the edge deletion methods, only if the perturbation level reaches 1.0, the ATS goes to 0, otherwise, the ATS remains higher compared to other settings. Relation replacement has a relatively stronger impact on ATS, i.e., on the semantics of the perturbed KGs. For SC2D and SD2 (which intuitively measure structural similarity), the four methods show similar levels of decrease in most cases.
Figure 6 illustrates the hallucination ratio using different methods on the perturbed KGs. We observe that (1) FS-SG consistently outperforms FS-T and FS-P even at higher perturbation levels, demonstrating its robustness by maintaining higher scores as perturbations increase. (2) FS-T and FS-P exhibit similar trends, each showing a significant performance drop as perturbation levels increase. Particularly, performance of FS-T and FS-P deteriorate when the perturbation level reaches 50%, i.e., becoming worse than the baseline using CoT. (3) On the setting using Relation Replacement which severely harms the semantics of the KGs, FS-T and FS-P decline more sharply than FS-SG. However, it still outperforms the baseline when the perturbation level is smaller than 40%.
Our findings on OKGQA-P illustrate that the effectiveness of KG-derived information diminishes with a perturbation level at 50%, surpassing this level leads to a further decrease in performance. We think that before this perturbation level at 50%, incorporating external knowledge from KGs can mitigate hallucinations in LLMs compared to baseline using CoT. In practical scenarios, platforms like Wikidata are less likely to be severely perturbed than 50% due to ongoing updates and community-based quality control, ensuring the relevance of our findings in real-world applications.
5 Conclusion
In this paper, we propose OKGQA and variant OKGQA-P, to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios. Unlike existing benchmarks that focus primarily on closed ended tasks, OKGQA presents diverse open-ended question types that mirror the unpredictable nature of practical applications. We conduct a series of experiments and analyze the effectiveness of various retrieval methods and LLMs of different magnitudes, providing insights for further research. Our results underscore the significance of integrating KGs with LLMs to help reduce hallucination of LLMs, even in circumstances where the KGs are contaminated.
Limitations
Our proposed benchmark primarily use DBpedia as the knowledge source, which may not generalize well to testing scenarios requiring highly specialized or domain-specific knowledge. Testing domain-specific open-ended QA may require constructing sub-graphs from domain-specific KGs. In addition, the study assumes a static KG for reasoning and retrieval. In dynamic environments where knowledge is continuously updated, maintaining and integrating real-time changes remains a challenge and may requires further design.
Appendix A Appendix
A.1 Intuition of Open-ended Questions?
In previous close-ended KGQA research Mihaylov et al. (2018); Yih et al. (2016); Talmor and Berant (2018), “hallucination” has not been considered at all. The primary reason is that close-ended KGQA models typically generate concise entity names as answers, which are directly compared against ground truth labels using metrics such as precision or accuracy. Any incorrect entity name is immediately reflected in the evaluation results, leaving little room for hallucination to manifest as a separate problem. While effective for simple tasks, we believe this closed-ended setting is too naive and fails to generalize to real-world, complex QA situations, where models must provide more detailed, nuanced responses Kantharaj et al. (2022).
In contrast, our work focuses on open-ended KGQA, where LLMs are prompted to generate more elaborate answers that include reasoning paths and supporting facts derived from the background KG. This broader output space not only provides a larger field to assess hallucination but also increases the likelihood of factual inaccuracies, as LLMs tend to produce more errors when generating longer and more complex responses. While hallucination is minimal in close-ended KGQA due to the short and factual nature of entity names, open-ended KGQA highlights the susceptibility of LLMs to hallucinate when generating rich, explanatory answers. By introducing this open-ended setting, OKGQA, we (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design. We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination.
A.2 Query Construction
In this section, we discuss the details of the query construction of OKGQA. It follows a template-based method with LLMs to generate a diverse range of queries. To ensure that the generated queries accurately represent real-world scenarios and complexities, we integrate the corresponding Wikipedia pages of each entity as contextual information for the LLMs’ generation. Furthermore, we follow a human-in-the-loop process to optimize the instruction used for generation, as shown in Figure 7. Intuitively, starting with an initial instruction, we generate a large number of query candidates, then use an LLM as an automatic evaluator to evaluate the quality of the query, denoted as a set of scores from different metrics $s_{\text{auto}}$ ranging from 1 to 10, with higher scores indicating better performance. Then, these queries are manually evaluated and assigned a set of human-label scores $s_{\text{human}}$ corresponding to $s_{\text{auto}}$, normalized to the same range as $s_{\text{auto}}$. We then optimize the input instructions by iteratively generating instructions to minimize the gap between $s_{\text{human}}$ and $s_{\text{auto}}$. This process quite mimics the way of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) and inherits the benefit that labeling the reward of the LLMs’ output is much easier than labeling the output directly.
Figure 7: Human-in-the-loop of query construction.
Specifically, we consider five angles to measure the quality of the generated queries: (1) Naturalness: assessing how fluid and human-like the query sounds; (2) Relevance: measuring whether the query pertains directly to the entity and the context provided; (3) Specificity: determining the level of detail and granularity included in the query, ensuring it is not too broad or vague; (4) Novelty: evaluating the uniqueness of the query, ensuring it is not just a repetitive or common question; (5) Actionability: gauging whether the query prompts clear, definite answers or actions that are feasible within the given context. Each of these angles contributes to a holistic evaluation of the query’s effectiveness and relevance in real-world applications.
We have three evaluators participating in the manual assessment of query quality. All of the evaluators are computer science majors with fluent English skills. As the evaluation centers on various linguistic metrics such as naturalness, relevance, specificity, novelty, and actionability, we only require the evaluators to possess a fundamental understanding of English without restricting their majors. We calculate the Pearson correlation coefficients between human and LLM scores as shown in Table 3. It shows that as the rounds progress, agreement between humans and LLMs increases, suggesting that iterative feedback improves alignment between human annotation and LLM responses.
Table 3: Pearson correlation coefficients between human and LLM scores across rounds.
Metric Round 1 Round 2 Round 3 Round 4
Naturalness 0.60 0.65 0.69 0.74
Relevance 0.55 0.59 0.64 0.70
Specificity 0.46 0.54 0.60 0.65
Novelty 0.49 0.57 0.63 0.67
Actionability 0.33 0.41 0.48 0.53
In addition, we also consider verifying the inter-rater reliability across three evaluators as shown in Table 4. We report the Cohen’s Kappa coefficient for each pair of evaluators as follows. Using the Landis and Koch (1977) interpretation guidelines, the Cohen’s Kappa coefficients for Naturalness and Relevance (ranging from 0.79 to 0.85) fall within the “Substantial” to “Almost Perfect” categories, indicating strong inter-rater reliability for these metrics. This reflects a shared understanding of the evaluation criteria, resulting in consistent ratings among evaluators. For Specificity, Novelty, and Actionability, the coefficients range from 0.58 to 0.68, placing them primarily in the “Moderate” to “Substantial” categories. These results suggest moderate reliability for these metrics, likely due to subjective interpretation and less clearly defined evaluation guidelines. Novelty, with lower coefficients around 0.61 to 0.63, highlights variability in ratings, suggesting that evaluators may have differing perspectives on what qualifies as novel (but the inter-rater reliability is still considered “Substantial”). Meanwhile, Actionability performs slightly better, nearing the “Substantial” range, indicating moderately consistent criteria.
Table 4: Cohen’s Kappa coefficient for various metrics.
Metric Evaluator 1 & 2 Evaluator 1 & 3 Evaluator 2 & 3
Naturalness 0.85 0.83 0.84
Relevance 0.81 0.79 0.80
Specificity 0.65 0.63 0.66
Novelty 0.60 0.58 0.61
Actionability 0.67 0.65 0.68
A.3 Personalized PageRank (PPR)
In this section, we discuss the details of the PPR algorithm used in §2.1 to prune the graph from DBPedia and concentrate on nodes most pertinent to the central nodes of interest. The PPR is calculated using the iterative formula:
$\mathbf{p}=\alpha\mathbf{A}^{\top}\mathbf{p}+(1-\alpha)\mathbf{s},$ (3)
[…]
A.4 Prize-Cost-based Path Retrieval
The path-retrieval method is designed to construct and evaluate paths in a graph based on predefined prize assignments and cost allocations. The objective is to form sequences of nodes and edges, represented as $\mathcal{P}=\{v_{1},e_{1},v_{2},\dots,e_{n-1},v_{n}\}$, that maximize the overall score and minimize the costs. To efficiently manage the exploration of potential paths, we utilize a priority queue, a data structure that allows paths to be organized based on their scores, ensuring that the highest-scoring paths are processed first. The method starts by picking a number of starting nodes with high prizes. We then expand these starting points by exploring neighboring nodes. For each neighbor, the method calculates a new score. This score is the sum of the neighbor’s prize and the edge’s prize minus the edge’s cost. If this neighbor hasn’t been visited before, which helps avoid looping, the algorithm adds this neighbor to the path. This new path is then added to the priority queue. This expansion keeps going until the path reaches a maximum length or can’t be extended further. The algorithm keeps track of paths already explored to avoid repetition and ensure paths don’t loop back on themselves. When no more paths can be added or the priority queue is empty, the algorithm sorts the paths by their scores from highest to lowest.
A.5 KG Similarity Metrics
To measure how much the perturbed KGs are deviated from the original KG used in §2. We re-present the according metrics from (Raman et al., 2020) below:
Aggregated Triple Score (ATS)  ATS measures semantic similarity between two KGs. Let $s_{\mathcal{G}}$ be an edge (triple) scoring function, such that $s_{\mathcal{G}}(e_{1},r,e_{2})$ measures how likely edge $(e_{1},r,e_{2})$ is to exist in $\mathcal{G}$. Also, assume $s_{\mathcal{G}}$ has been pre-trained on $\mathcal{G}$ for link prediction. Then, ATS is defined as $f_{\text{ATS}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=\frac{1}{|\mathcal{T}^{{}^{\prime}}|}\sum_{(e_{1},r,e_{2})\in\mathcal{T}^{{}^{\prime}}}s_{\mathcal{G}}(e_{1},r,e_{2})\in[0,1]$, which denotes the mean $s_{\mathcal{G}}$ score across all edges in $\mathcal{G^{\prime}}$. Intuitively, if a high percentage of edges in $\mathcal{G^{\prime}}$ are also likely to exist in $\mathcal{G}$ (i.e., high ATS), then we say that $\mathcal{G^{\prime}}$ and $\mathcal{G}$ have high semantic similarity. $s_{\mathcal{G}}$ is task-specific, as KGs from different tasks may differ greatly in semantics. We use the $s_{\mathcal{G}}$ from (Li et al., 2016); while ATS captures semantic KG differences, it is not sensitive to KG connectivity structure. Note that $f_{\text{ATS}}(\mathcal{G},\mathcal{G})$ may not equal 1, since $s_{\mathcal{G}}$ may not perfectly generalize to KGs beyond those it was trained on.
Similarity in Clustering Coefficient Distribution (SC2D)  SC2D measures structural similarity between two KGs and is derived from the local clustering coefficient (Saramäki et al., 2007; Onnela et al., 2005; Fagiolo, 2007). For a given entity in $\mathcal{G}$ (treated here as undirected), the local clustering coefficient is the fraction of possible triangles through the entity that exist (i.e., how tightly the entity’s neighbors cluster around it). For entity $e_{i}\in\mathcal{E}$, the local clustering coefficient is defined as $c_{i}=2\text{Tri}(e_{i})/(\text{deg}(e_{i})(\text{deg}(e_{i})-1))$, where $\text{Tri}(e_{i})$ is the number of triangles through $e_{i}$, and $\text{deg}(e_{i})$ is the degree of $e_{i}$. For each relation $r\in\mathcal{R}$, let $\mathcal{G}^{r}$ be the subgraph of $\mathcal{G}$ consisting of all edges in $\mathcal{T}$ with $r$ That is, $\mathcal{G}^{r}=(\mathcal{E},r,\mathcal{T}^{{}^{\prime}})$, where $\mathcal{T}^{{}^{\prime}}=\{(e,r,e^{\prime})\hskip 2.84526pt|\hskip 2.84526pte,e^{\prime}\in\mathcal{E}\}$. Let $\mathbf{c}^{r}$ denote the $|\mathcal{E}|$-dimensional clustering coefficient vector for $\mathcal{G}^{r}$, where the $i$th element of $\mathbf{c}^{r}$ is $c_{i}$. Then, the mean clustering coefficient vectors for $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are $\mathbf{c}_{o}=\frac{1}{|\mathcal{R}|}\sum_{r\in\mathcal{R}}\mathbf{c}^{r}$ and $\mathbf{c}_{p}=\frac{1}{|\mathcal{R}^{\prime}|}\sum_{r\in\mathcal{R}^{\prime}}\mathbf{c}^{r}$, respectively. SC2D is defined as $f_{\text{SC2D}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{c}_{o}-\mathbf{c}_{p}\rVert_{2}}{\lVert\mathbf{c}_{o}-\mathbf{c}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.
Similarity in Degree Distribution (SD2)  SD2 also measures structural similarity between two KGs, while addressing SC2D’s ineffectiveness when the KGs’ entities have tiny local clustering coefficients (e.g., the item KG used by recommender systems is roughly bipartite). In such cases, SC2D is always close to one regardless of the perturbation method, thus rendering SC2D useless. Let $\mathbf{d}^{r}$ denote the $|\mathcal{E}|$-dimensional degree vector for $\mathcal{G}^{r}$, where the $i$th element of $\mathbf{d}^{r}$ is $\text{deg}(e_{i})$. Then, the mean degree vectors for $\mathcal{G}$ and $\mathcal{G}^{\prime}$ are $\mathbf{d}_{o}=\frac{1}{|\mathcal{R}|}\sum_{r\in\mathcal{R}}\mathbf{d}^{r}$ and $\mathbf{d}_{p}=\frac{1}{|\mathcal{R}^{\prime}|}\sum_{r\in\mathcal{R}^{\prime}}\mathbf{d}^{r}$, respectively. SD2 is defined as $f_{\text{SD2}}(\mathcal{G},\mathcal{G}^{{}^{\prime}})=1-\frac{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}}{\lVert\mathbf{d}_{o}-\mathbf{d}_{p}\rVert_{2}+1}\in[0,1]$, with higher value indicating higher similarity.
Appendix B Related Work
Due to the stochastic decoding process of Large Language Models (LLMs), i.e., sampling the next token in the sequence, LLMs exhibit probabilistic behaviors: (1) potentially yielding varied outputs of the same input across different instances (Agrawal et al., 2023); (2) cannot accurately interpret phrases or terms when the context is vague and resides in a knowledge gap of the model. This will lead to outputs that may sound plausible but are often irrelevant or incorrect. This will lead to outputs that may sound plausible but are often irrelevant or incorrect. This “hallucinations” undermines the reliability of LLMs (Huang et al., 2023). One emerging research trend is enhancing LLMs through integrating external knowledge graphs (Agrawal et al., 2023). KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning (Zheng et al., 2023; Agrawal et al., 2023; Sui et al., 2022). Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent.
Researchers employ diverse strategies to augment the LLMs by integrating external KGs (Sui et al., 2024; He et al., 2024b). For example, KAPING (Baek et al., 2023b) matches entities in questions to retrieve related triples from knowledge graphs for zero-shot question answering. Wu et al. (2023) finds that converting these triples into textualized statements can further enhance LLM performance. StructGPT (Jiang et al., 2023b) propose to convert user query into structured formats (e.g., SPARQL) for information extraction from KGs. Following the succuess of internal reasoning-enhancement methods like Chain-of-thoughts (CoT) (Wei et al., 2022), Reflexion (Shinn et al., 2024), and Tree-of-thoughts (ToT), He et al. (2022) propose “rethinking with retrieval” to use decomposed reasoning steps from CoT prompting to retrieve external knowledge, leading to more accurate and faithful explanations. IR-CoT (Trivedi et al., 2022) interleaves the generation of CoT with knowledge retrieval from corresponding KGs, iteratively guiding both retrieval and reasoning for multi-step questions. MindMap (Wen et al., 2023) introduce a plug-and-play approach to evoke graph-of-thoughts reasoning in LLMs. Similarly, RoG (Luo et al., 2023b) use KGs to create faithful reasoning paths based on various relations, enabling interpretable reasoning in LLMs.
However, current benchmarks for testing the capabilities of these LLM+KG models are predominantly closed-ended, restricting responses to a limited set of entities/relations or a set of logical forms derived from specific facts of KG. Hence, they can only test a very limited subset of the LLM’s tendency to hallucinate, leaving a gap in the assessment of complex, real-world scenarios. Particularly, standard metrics such as FActScore (Min et al., 2023) and SAFE (Wei et al., 2024) for evaluating the hallucination rate of LLMs require open-ended settings, i.e., questions are phrased as a statement which requires a longer answer. Compared with previous works, our proposed OKGQA is tailored for evaluating LLMs enhanced with KGs under open-ended, real-world question-answering scenarios. The benchmark extends the assessment of closed-ended question answering to an open-ended setting, which can further support the assessment of hallucination of LLMs.
Appendix C Prompt List
In this section, we present all the prompts required for the main experiments. To enhance clarity, we provide only one example in the prompt labeled as Example 1; the other few-shot examples utilized are labeled as Other In-Context Few-shots within the prompt.
C.1 Knowledge-Augmented Generation
System Instruction: “You are a helpful assistant designed to answer the users’ open-ended questions. Your task is to provide accurate, concise, and useful information to foster understanding and solve problems. Whether the questions relate to complex scientific concepts, historical events, practical advice, or everyday life, your goal is to assist by offering thoughtful and informative responses.”
In-Context Few-shots
Question: {Question}
Below are the facts that might be relevant to answer the question: {Retrieved-knowledge}
Answer:
C.2 OKGQA Query Generation Prompting
System Instruction: “Generate open-ended questions about different types: character description, event description, cause explanation, relationship explanation, trend prediction, outcome prediction, contrast analysis, historical comparison, evaluation and reflection, and application and practice. Please provide specific suggestions. Generate the questions, the type of the questions, the placeholders, the naturalness of your generated questions (choose from high, medium, and unnatural), the difficulty of the generated questions (choose from hard, medium and easy) and DBPedia_entities (link the placeholders to DBPedia entities) in JSON format.”
Example 1: as shown in Figure 8.
Figure 8: Example 1 Demonstration.
Other In-Context Few-shots
Generation:
C.3 Prompts for Instruction Tuner
Act as an “Instruction Tuner” for the LLM, you will be given the inputs: (1) the {Current Instruction} used to guide the LLMs’s evaluation, including specific examples with ground truth labels; (2) {Current Errors} that emerged with this instruction are applied to the dataset.
The current errors are presented in the following format: (1) INPUT: {input text} (2) PREDICTED OUTPUT: {predicted label}, (3) EXPECTED OUTPUT: {ground truth label}. Carefully analyze these errors and craft a revised concise instruction for the LLM to fit the expected outputs. Include 2-3 examples at the end of your response to demonstrate how the new instruction would be applied.
C.4 Metrics Prompt for G-eval
System Instruction: “You are a helpful assistant designed to evaluate the quality of the response to a query. Your task is to rate the response on one metric defined as below:”
Empowerment Criteria: Evaluate whether the “Actual Output” can help the reader understand the topic and make informed decisions regarding the “Input”. A response with high empowerment provides accurate information and explanations that enhance the reader’s understanding. When evaluating empowerment, consider the relevance of the information provided in the “Actual Output” to the “Input” and the “Retrieval Context”.
Comprehensiveness Criteria: Evaluate the extent to which the “Actual Output” covers all aspects and details of the question “Input”. A comprehensive answer should thoroughly address every part of the question, leaving no important points unaddressed. When evaluating comprehensiveness, consider the relevance of the information provided in the “Actual Output” to the “Input” and the “Retrieval Context”.
Correctness Criteria: Measure how clearly and specifically the “Actual output” responds to the question “input”. A highly direct response stays focused on the question, providing clear and unambiguous information. When evaluating correctness, consider the relevance of the information provided in the “Actual Output” to the “Input” and the “Retrieval Context”.
Context Relevance Criteria: Evaluate the extent to which the “Actual output” incorporates relevant information from the “Retrieval Context”. This includes assessing whether the output adheres to the thematic, factual, and situational specifics presented in the “Retrieval Context”. Relevant responses not only address the direct query but also align closely with the contextual elements provided, ensuring a seamless and coherent transition between the “Retrieval Context” and the “Actual Output”. The most contextually relevant responses demonstrate an understanding and appropriate reflection of the given circumstances, historical facts, or conceptual background, thereby contributing to the overall accuracy and utility of the information provided.
Response: [Respond with metric and the corresponding score.]